OpenStudy (anonymous):

: Solve 3/(4q)+1/(2q^2)=1/(8q). i got q=2/3…

OpenStudy (anonymous):

q would actually be -4/5. do you want me to type out the steps?

OpenStudy (anonymous):

can you draw it

OpenStudy (anonymous):

\[1/(2 q^2)+3/(4 q) = 1/(8 q)\] Subtract 1/(8 q) from both sides: \[1/(2 q^2)+5/(8 q) = 0\] Multiply both sides by q^2: \[(5 q)/8+1/2 = 0\] Subtract 1/2 from both sides: \[(5 q)/8 = -1/2\] Divide both sides by 5/8: \[q = -4/5\]

OpenStudy (anonymous):

is this a different way to solving rational equations

OpenStudy (anonymous):

No i think this would be the only way to do it. Or at least the most logical...

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no prob